loggies room has a perimeter of 120 meters. the room is the shape of a rectangle which nthe width is 5 times the length. find the width of loggies room.
Answer by checkley79(3341)

P=2L+2W
W=5L
120=2L+2*5L
120=2L+10L
120=12L
L=120/12
L=10 METERS.
W=5*10=50 METERS.
PROOF:
120=2*10+2*50
120=20+100
120=120